in general form find the equation of the line through (-4,0) and (3,-5)
Answer by stanbon(75887)

in general form find the equation of the line
through (-4,0) and (3,-5)
-----
slope = (0--5)/(-4-3) = 5/-7
----
Form y = mx+b; solve for "b".
---
0 = (-5/7)(-4) + b
0 = 20/7 + b
b = -20/7
-------
Equation:
y = (-5/7)x - (20/7)
---
Standard form:
5x + 7y = -20
=================
Cheers,
Stan H.
=================